OpenStudy (anonymous):

hw do u solve square root over 6-p=p?

OpenStudy (anonymous):

Do you mean \[\sqrt{p}/(6-p) = p\]

OpenStudy (anonymous):

nevermind, you made a different question

OpenStudy (radar):

\[\sqrt{6-p}=p\] square both sides getting:\[6-p=p ^{2}\]subtract 6-p from both sides getting: \[p ^{2}+p-6=0\] factor and set to 0 (p+3)(p-2)=0 solve for p: p=-3, p=2